find the slope: 36x – 9y = 15
Answer by CharlesG2(834)

find the slope: 36x – 9y = 15

okay so we need to solve for y
36x - 9y = 15
-9y = -36x + 15 (rearranged terms)
9y = 36x - 15 (divided both sides by -1)
y = (36/9)x - (15/9) (divided both sides by 9)
y = 4x - 5/3 (simplified)
now it is in y = mx + b form, where m is the slope,
and b is the y-intercept (vertical-intercept or point (0,b))

slope = m = 4